Please help me with this question: Find all real zeros of the polynomial function: f(x)=-6x^4+54x^2
Answer by lwsshak3(11628)

Find all real zeros of the polynomial function:
f(x)=-6x^4+54x^2
set f(x)=0
-6x^4+54x^2=0
6x^4-54x^2=0
x^2(6x^2-54)=0
x^2=0
x=0
or
6x^2-54=0
6x^2=54
x^2=54/6=0
x=±√9=±3
zeros are: 0, 0, -3, 3